# Large Language Model for Causal Decision Making

### 29 Dec 2023 v2

@@ Abstract
Large Language Models (LLMs) have shown their success in language understanding and reasoning on general topics. However, their capability to inference based on user-specified structured data and knowledge in corpus-rare concepts like causal decision-making is still limited. In this work, we explore the possibility of fine-tuning an open-sourced LLM into LLM4Causal, which can identify the causal task, execute a corresponding function, and interpret its numerical results based on users’ queries and the provided dataset. Meanwhile, we propose a data generation process for more controllable GPT prompting and present two instruction-tuning datasets: (1) Causal-Retrieval-Bench for causal problem identification and input parameter extraction for causal function calling and (2) Causal-Interpret-Bench for in-context causal interpretation. With three case studies, we showed that LLM4Causal can deliver end-to-end solutions for causal problems and provide easy-to-understand answers. Numerical studies also reveal that it has a remarkable ability to identify the correct causal task given a query. 00footnotetext: *Corresponding author: Rui Song, songray@gmail.com. This work is not related to the author’s position in Amazon. 00footnotetext: +Equal Contribution
@@ 1 Introduction
Recently, Transformer-based LLMs containing billions of parameters are gaining popularity and are widely applied in fields such as education, legal services, and medicine (Kasneci et al., 2023; Chen et al., 2021). LLMs such as GPT-3 (Brown et al., 2020a), GPT-4, and LLaMA (Touvron et al., 2023) have shown impressive performance in multiple natural language processing tasks such as question-answering, common-sense reasoning, and translation (Zhao et al., 2023). These superior performances of LLMs have also motivated explorations on LLMs’ application on the causal decision-making process.
To explore LLMs’ potential for causal decision-making processes, recent studies are majorly based on LLM’s knowledge saved internally in its parameters. For example, Kasneci et al. (2023) utilizes LLM to recover pairwise causal relationships based on the metadata of the variables, instead of the data values. Long et al. (2023) further extends the approach to obtain causal graph estimates by querying LLM about the variable pairs. However, this line of approach has some limitations: First, this approach is limited to causal structure learning and cannot be applied to other causal tasks such as causal effect estimation and causal policy learning. Second, this type of approach is heavily based on internal knowledge in the LLM training corpus, which can be incomplete and outdated. Third, this approach lacks step-by-step reasoning and provides the results in a black-box way.
A natural candidate to overcome the above limitations is to ask the LLM to complete the causal tasks utilizing data input. Two notable examples of such data-enhanced LLMs include i) the Advanced Data Analysis plugin of ChatGPT (OpenAI, 2023a), formerly known as the Code Interpreter (OpenAI, 2023b), which allows users to instruct the LLM through prompts to conduct technical data analysis on self-uploaded datasets; and ii) the latest GPT-3.5/GPT-4 models that include the function calling feature (OpenAI, 2023c), allowing users to introduce new functions for LLM to select from, by providing detailed descriptions and structured information about the function. Additionally, many other recent developments like ToolLLM(Qin et al., 2023b) and GPT4Tools(Yang et al., 2023) also enable LLMs to use external tools to conduct data analysis but are not specifically tailored for the causal tasks.
Despite their ability to analyze external datasets, all of the aforementioned data-enhanced LLMs have the following two limitations in dealing with causal problems: First, in-depth domain knowledge is required to validate the results provided by the current GPT-based models. While these models can provide a wealth of information related to the user’s query, they frequently hallucinate and mislead the user with irrelevant context. For example, they may provide correlation analysis results when causal effect analysis is expected, as shown in Figure 1(a). Second, while some existing approaches may provide guidance for implementing some classical causal methods and suggest the corresponding code script, they fail to provide end-to-end result delivery and easy-to-understand interpretation. Third, pre-trained on massive text corpus, it only provides executions for major methods contained in the training data and requires users’ effort for implementing newly-released approaches as shown in Figure 1(b).
@@
To overcome these challenges, in this study, we develop an end-to-end user-friendly large language model with causal decision-making ability for general audiences. As illustrated in Figure 1, the proposed LLM4Causal model consists of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation. Given a user query and a data file uploaded by the user, the first step transforms the query in natural language to a structured summary in JSON format, which contains essential information such as dataset name, task type, and variable of interest. This transformation is conducted by fine-tuning a pre-trained LLM on the synthetic data, where a well-designed data generation pipeline, is proposed to improve the data quality in terms of both data variety and accuracy. Given the structured JSON data, the second step automatically selects the causal learning algorithm based on the task type and executes it to generate model outputs. The model outputs are then converted to easy-to-understand languages via the third step, where the same LLM model is further fine-tuned on additional data to produce interpretations adhering to three commonly used evaluation rubrics.
@@
Figure 1: A flowchart of the proposed LLM4Causal model consisting of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation.
@@
  • This paper is the first attempt to construct an end-to-end user-friendly large language model, LLM4Causal, with causal decision-making ability, to the best of our knowledge. Without requiring in-depth domain knowledge, the LLM4Causal could be easily used for general audiences, which addresses the weakness of the current LLM applications on causal tasks.
  • We have finetuned the LLM4Causal to equip it with the capability to conduct end-to-end analysis on causal-related tasks in three steps: i) interpreting user request by converting it to a structured data format, ii) assigning causal tools and executing the corresponding algorithm, and iii) providing easy-to-understand interpretation on the algorithm output.
@@
  • We have proposed a three-step data generation pipeline that combines LLM text generation and human annotation to create customized input-output data. With this well-designed pipeline, we have collected two high-quality benchmark datasets, Causal-Retrieval-Bench for causal function calling and Causal-Interpret-Bench for causal interpretation, with outstanding data variety and accuracy.

### 12 Apr 2024 v3

# LLM4Causal: Large Language Model for Causal Decision Making

@@ Abstract
Large Language Models (LLMs) have shown their success in language understanding and reasoning on general topics. However, their capability to perform inference based on user-specified structured data and knowledge in corpus-rare concepts, such as causal decision-making is still limited. In this work, we explore the possibility of fine-tuning an open-sourced LLM into LLM4Causal, which can identify the causal task, execute a corresponding function, and interpret its numerical results based on users’ queries and the provided dataset. Meanwhile, we propose a data generation process for more controllable GPT prompting and present two instruction-tuning datasets: (1) Causal-Retrieval-Bench for causal problem identification and input parameter extraction for causal function calling and (2) Causal-Interpret-Bench for in-context causal interpretation. By conducting end-to-end evaluations and two ablation studies, we showed that LLM4Causal can deliver end-to-end solutions for causal problems and provide easy-to-understand answers, which significantly outperforms the baselines.
@@ 1 Introduction
Recently, Transformer-based LLMs containing billions of parameters are gaining popularity and are widely applied in fields such as education, legal services, and medicine (Kasneci et al., 2023; Chen et al., 2021). LLMs such as GPT-3 (Brown et al., 2020), GPT-4, and LLaMA (Touvron et al., 2023a) have shown impressive performance in multiple natural language processing tasks such as question-answering, common-sense reasoning, and translation (Zhao et al., 2023). These superior performances of LLMs have also motivated explorations on LLMs’ applications on the causal decision-making (Song et al., 2023) procedures, which include various crucial tasks in real-life such as causal structure learning (CSL) tasks (Spirtes et al., 2000; Glymour et al., 2019), causal effect learning (CEL) tasks (Yao et al., 2021; Hicks and Tingley, 2011) and causal policy learning (CPL) tasks (Chakraborty and Murphy, 2014; Sutton and Barto, 2018; Zeng et al., 2023).
Most of recent works that applied LLM in causal decision-making tasks are focused on CSL, where LLM’s internal knowledge gained during its training process is exploited in learning the causal relations among variables. For instance, both Kıcıman et al. (2023) and Long et al. (2023) propose to obtain the causal relationships between variables by directly querying GPT models using the variables names, assuming that LLMs can extract internal causal knowledge of those variables from the large corpus it was trained on. In a further development by Ban et al. (2023), LLM’s internal causal knowledge are utilized to guide the optimization process of score-based methods as soft or hard constraints, enhancing the performance compared to using the traditional CSL methods alone. However, this line of works heavily rely on the quality of the knowledge in LLM’s training corpus and lacks step-by-step reasoning or interpretation. A natural choice to overcome these limitations is to ask the LLM to complete the causal tasks utilizing existing causal decision making tools (such as CausalML(Chen et al., 2020), CausalDM (Song et al., 2023), causal-learn (Zheng et al., 2023)) and input data uploaded by users. In literature, Code Interpreter (OpenAI, 2023a), ToolLLM (Qin et al., 2023b) and GPT4Tools(Yang et al., 2023) have leveraged the function calling feature of LLM in general data analysis tasks for user input datasets. However, directly applying those methods on causal related tasks and datasets may lead to several issues: first, since these models are not tailored on causal tasks, they frequently hallucinate and mislead the user with irrelevant context, such as providing correlation analysis results when causal effect analysis is expected, as shown in Figure 0(a); second, most of these methods fail to provide end-to-end result delivery (see Figure 0(b)) and easy-to-understand interpretation for causal related tasks; third, these methods lack information about newly released approaches that are not included in LLM’s training corpus (see Figure 0(c)).
(a) User queries on causal relationships.
(b) User queries on a classical method, LiNGAM(Shimizu, 2014).
(c) User queries on a newly-developed method, ANOCE (Cai et al., 2020).
Figure 1: User interaction with ChatGPT on causal related questions. [^1] The chat histories could be found in link 1, link 2, and link 3.
@@
To overcome these challenges, in this study, we develop an end-to-end user-friendly large language model with causal decision-making ability for general audiences. As illustrated in Figure 2, the proposed model consists of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation. Upon receiving a user query and an uploaded data file, the initial step identifies the pertinent causal task and extracts query details—including dataset name, task type, and variable of interest, among others—into a structured JSON summary. The LLM4Causal model gains the ability to convert natural language user queries into JSON summaries by fine-tuning a pre-trained LLM on the Causal-Retrieval-Bench dataset. A well-designed data generation pipeline is proposed to ensure the quality of Causal-Retrieval-Bench, in terms of both data variety and accuracy. In the second step, the system automatically chooses the causal learning algorithm according to the task type detailed in the structured JSON data, executes the selected algorithm to analyze the dataset, and collects the algorithm’s output. The outputs are then translated into easily understandable language in the third step, using the LLM4Causal model, which has been further fine-tuned with the Causal-Interpret-Bench dataset to generate high-quality interpretations.
@@
Figure 2: A flowchart of the LLM4Causal consists of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation.
@@
  • This paper is the first to construct an end-to-end user-friendly large language model (LLM4Causal) with causal decision-making ability. LLM4Causal could be easily used for general audiences, which addresses the weakness of the current LLM applications on causal tasks. It has the capability of i) interpreting user requests by causal task classification and information extraction, ii) assigning causal tools and executing the corresponding algorithm, and iii) providing an easy-to-understand interpretation of the algorithm output.
@@
  • We have proposed a three-step data generation pipeline that combines LLM text generation and human annotation to create fine-tuning datasets. With this well-designed pipeline, we have collected two high-quality benchmark datasets, Causal-Retrieval-Bench for causal function calling and Causal-Interpret-Bench for causal interpretation, with outstanding data variety and accuracy.